Lakersfan23:

Identify the expression as an equation containing fractions, a fractional equation, or a proportion. x/12=17/3-2x/9-5+3x/12

jhonyy9:

\[\frac{ x }{ 12 } = \frac{ 17 }{ 3 }-\frac{ 2x }{ 9 } + 5 + \frac{ 3x }{ 12 }\] do you think on this in this way ?